OpenStudy (anonymous):

A die is rolled 12 times. Find the chance that the face with six spots appears once among the first 6 rolls, and once among the next 6 rolls.

OpenStudy (kropot72):

The binomial distribution applies here: \[P(once\ among\ 6\ rolls)=\left(\begin{matrix}6 \\ 1\end{matrix}\right)(\frac{1}{6})(\frac{5}{6})^{5}=(\frac{5}{6})^{5}=you\ can\ calculate\]

OpenStudy (anonymous):

0.1615